# https://www.intelligence.tuc.gr/gogs/

### v2

@@ 3.1 Multiagent learning
Another notable paper is [chu2019multi] that introduces a fully scalable and decentralized MADRL algorithm, the Multiagent Advantage Actor Critic (MA2C). MA2C is built as an extension to the well-known DRL algorithm Advantage Actor Critic (A2C) and is designed to tackle the adaptive traffic signal control problem. It does so, by enabling limited communication between the agents, and specifically sharing fingerprints of policies between neighbors. The connections between the agents help to alleviate the issues caused by partial observability and non-stationarity, as it helps each entity to have more informed decision-making based on the anticipated actions of its neighbors. Another key feature of this framework, is that it tailors the reward of each agent $i$, by computing a discounting, weighted, global reward with increased weights for the rewards of agents in agent’s $i$ neighborhood. Essentially, each individual’s reward is discounted according to the agent’s distance from agent ii. MA2C is evaluated on both a large synthetic traffic grid and the real-world traffic network of Monaco city. In this context, the algorithm demonstrates superior performance of optimality, robustness and sample efficiency over traditional decentralized MADRL algorithms and other state-of-the-art methods.
@@
Despite the fact that they can be criticized as lacking in theoretical foundations in a multiagent setting, independent learners (IL) can attain high performance in hard multiagent problems. The underpinning of IL algorithms is that each agent learns on its own and perceives the rest of the agents as part of the environment with no communication in between, thus pertaining to the SIH. This simple logic can surprisingly attain performance comparable, and sometimes even better than more sophisticated techniques [de2020independent, papoudakis2020benchmarking, yu2022surprising], and simultaneously be scalable due to its by-design decentralization. The first algorithm introduced in this context is a value-based algorithm based on $Q$-learning, the Independent $Q$-learners (IQL) [tan1993multi]. According to IQL, each agent maximizes their own $Q$-function independently and optimize their own policy without sharing any information and without cooperating. A more modern approach adapts PPO in independent multiagent environments to present the actor-critic algorithm called Independent Proximal Policy Optimization (IPPO) [de2020independent]. In a similar fashion to IQL, IPPO does not require any information sharing or communication; instead, each agent works towards optimizing its own policy. Despite that, IPPO has achieved very good performance in hard problems such as the Starcraft Multiagent Challenge. Although IL can achieve competitive results and at the same time be straightforward and simplistic, it may struggle with the non-stationarity of the environment that is introduced by other agents’ behavior, since each decision-maker acknowledges its equivalents as dynamic parts of the environment.
@@
A relevant algorithm that uses a more sophisticated approach and is based on PPO, is that of Multiagent Proximal Policy Optimization (MAPPO) [yu2022surprising] that achieves good results in several cooperative benchmarks. MAPPO is an actor-critic algorithm, that follows a centralized training architecture which in turn leverages global information as the critic is trained on the joint state value function, making the agents more altruistic and disregarding the SIH. Additionally, there is no communication incorporated in this technique since the agents train in a centralized manner. The result is a sample-efficient algorithm that, unlike IL approaches, mitigates the non-stationarity problem making it less susceptible to changes in the environment made by other learners. As its single-agent counterpart, MAPPO serves as a very strong baseline due to its high performance in most cooperative environments and its robustness to hyperparameter tuning [papoudakis2020benchmarking].
@@
Another interesting approach is the Value Decomposition Networks (VDN) [sunehag2017value]. VDN is a groundbreaking algorithm that is designed to train multiple agents to work together towards achieving a common goal. Its key point, as the name suggests, is the principle of decomposing the joint value function into a sum of individual agents’ value functions. Ultimately, each agent builds a policy that maximizes the sum of policies which serves as a proxy for the joint value function. This offers a subtle way of coordination without the use of explicit communication mechanisms, since each entity makes decisions based on its perspective while indirectly contributing to collective intelligence. By utilizing the sum of these individual contributions and matching local rewards to the corresponding agents, VDN tackles two inherent MAS pathogenies such as the credit assignment problem and scalability. This method has opened up new avenues in research for designing sophisticated MAS capable of complex interactions and deeply strategic behavior.
@@
Finally, the Cooperative Multi-Agent Exploration (CMAE) [liu2021cooperative] is an exploration strategy that works on top of any MARL algorithm and promotes coordination through shared goals, particularly in sparse-reward environments. CMAE innovates by projecting the complex, high-dimensional state space into simpler, lower-dimensional “restricted spaces” for more efficient exploration. Using a hierarchical approach, exploration starts from these simpler spaces and expands as needed. This mechanism reshapes the rewards within the replay buffer to incentivize agents towards common objectives. It relies on entropy-based techniques for goal selection and demonstrates significant improvements in learning efficiency and effectiveness in tasks within environments like the Multiple Particle Environment and the StarCraft Multi-Agent Challenge, showcasing its superiority over traditional exploration methods.
@@
MA2C [chu2019multi] ATOC [jiang2018learning] VDN [sunehag2017value] MAPPO [yu2022surprising] IQL [tan1993multi] IPPO [de2020independent]
@@ 4 Rationality Solution Concepts
In complex multi-agent domains, understanding Nash equilibrium (NE) and other rationality solution concepts is vital for designing agents that interact effectively, as well as analyzing and providing insights on their behavior and outcomes. Usually, the desired properties of the outcomes in this scope are fairness and stability. Fairness corresponds to how well the agents’ payoffs comply with their actual contributions, while stability reflects the non-willingness of agents to select a deviating strategy [chalkiadakis2022computational]. In this section, we will present and discuss such game theoretic notions within the field of multi-agent DRL, with an emphasis on correlated equilibrium (CE), and on fairness-related solution concepts.
To begin with, the well-known Nash theorem guarantees that every finite two-player game has at least one NE. In games where multiple equilibria exist, one has to search for the most desirable one among them. One measure of desirability is the degree of social welfare maximization, i.e. the sum of all players’ payoffs. The search for an NE is PPAD-complete, however finding an NE with maximal social welfare is NP-hard even in symmetric 2-player games, making it computationally challenging. There have been attempts to reach approximate Nash equilibria that achieve near optimal social welfare. In [czumaj2015approximate] it is shown that this can be discovered in polynomial time.
@@
Moreover, we feature some successful integrations of CE in multi-agent DRL. As an example, in [yu2014multi] the authors apply CE on smart generation control in power systems. They present a novel algorithm that learns to achieve optimal joint equilibrium strategies. This approach employs the correlated equilibria in order to facilitate optimal coordination and consequently improve performance. Similarly, in [gan2019new] they employ a dynamic correlation matrix to model agent interactions and complex relationships. With this technique, the authors enable the discovery of optimal CE and enhance collective outcomes. This work highlights how valuable are CE in hostile environments were agent interconnections are complex and change dynamically. Another successful application is presented in [he2022multi], where multiple agents employ the DQN algorithm to achieve CE optimal solutions in order to optimize a textile manufacturing optimization process.
In [graf2024computational], authors employ DRL for training agents that participate into uniform price auctions. Interestingly, this Deep Deterministic Policy Gradient (DDPG) algorithm leads to learned policies which are shown to converge to NE. Also, this algorithm can tackle continuous state and action spaces, a feature that Q-learning does not possess, though a drawback of this approach is that it requires fine tuning of the related parameters.
@@
A notable work that addresses fair solutions is [jiang2018learning]. In this work, authors present a novel Fair-Efficient Network (FEN) model that manages to learn both efficiency and fairness in multiagent domains. This is achieved via fairness decomposition into objectives and utilizes a hierarchical structure, which contains a controller and multiple sub-entities, each one maintaining a distinct policy. This method supports fairness and addresses the challenges of stability in learning environments, as it can scale to larger problems due to the model’s ability to facilitate decentralized learning. Another notable contribution, is that of [zimmer2021learning], where a metric is put forward that evaluates fairness in the form of optimizable welfare functions. A novel NN architecture is used that consists of sub-networks and is capable of optimizing the different aspects of fairness such as efficiency and equity in the form of welfare functions. Ultimately, this technique leads to fair outcomes. Additionally, in [wang2020multi], a novel framework for controlling Unmanned Aerial Vehicles (UAVs) and optimizing their trajectories in terms of both performance and fairness is presented. In this context, authors propose a method where each UAV maximizes its reward independently and, at the same time, a collective approach for maximizing fairness is maintained.
@@
Finally, we list the most common solution concepts that can be found in the MARL literature. These are the Shapley value [shapley1953value], the Banzhaf index [banzhaf1964weighted], the core [gillies1959solutions], the nucleolus [schmeidler1969nucleolus], the kernel [davis1965kernel], the bargaining set [aumann1964bargaining], and the stable set [von1947theory].
The Shapley value guarantees that an agent will receive a payoff that is proportional to its contribution. Considering a characteristic function game $G=(N,v)$ with NN agents and $v$ the characteristic function, then the Shapley value $\phi_{i}(G)$ is calculated based on each agent’s $i$ marginal contribution:
@@
While the Shapley value and the Banzhaf index focus on the fairness of the payoffs, there are also solution concepts that deal with the stability of the formed coalitions in the game. The core solution concept is the set of all outcomes $(CS,\mathbf{x})$, where CSCS is a coalition structure and $\mathbf{x}$ the payoff vector, for which $x(C)\geq v(C)$ for every $C\subseteq N$, where $x(C)$ denotes the total payoff for $C$ under $\mathbf{x}$. Note that not every game has a non-empty core.
@@
The works of [chalkiadakis2007bayesian, 10.1145/1329125.1329203] combined the concept of Bayesian core, a cooperative solution concept, with non-cooperative solution concepts. This is inline with the our objectives, since we envision to revolutionize the way we think about equilibria, and pave the road towards the discovery of novel, meaningful, and practical concepts of rationality, so as to enable computationally tractable strategic deliberations in uncertain multiagent environments. Recall that a grand challenge put forward early in this paper, is that of overriding restricting GT assumptions, such as the SIH and CPA, [^1] Note that the Bayesian core solution concept does not necessitate the use of the CPA assumption. and to identify tractable ways for computing initial heterogeneous agent beliefs into the focus. This could potentially give rise to novel, practical equilibrium concepts, which explicitly take the initial agent beliefs-forming stage into account - and which come complete with bounded rationality capabilities and anytime performance guarantees. Such equilibrium concepts would have the potential to be employed in cooperative and non-cooperative domains alike, rendering as such the distinction between cooperative and non-cooperative GT solution concepts artificial and obsolete.
@@ 5 Probabilistic Topic Modeling
Existing Probabilistic Topic Modeling (PTM) algorithms have been applied successfully in various domains, such as document and text classification, image analysis and organization, movie recommendation systems, and population genetics [blei2012]. One could potentially adapt existing PTM algorithms for opponent modeling or strategy modeling purposes in multiagent environments—i.e., use them to derive the prior beliefs of the agents. Note that the ability to assign meaningful probability to previously unseen “documents” is inherent in many existing PTM algorithms (such as the LDA). Note also that PTMs constitute Bayesian exploration techniques, assigning as they do “topics” to modeled items according to probabilistic beliefs embodied in the mixture distributions they maintain. As such, they do not require the use of explicit heuristics or stoppage conditions to tackle the exploration-exploitation trade-off that is ever-present in decision-making problems.
Probabilistic topic models (PTMs) utilize statistical methods to analyze words in documents, so as to discover the underlying topics and their interconnection. The most popular implementation is based on the Latent Dirichlet Allocation (LDA) [blei2003].
@@
However, computing the probability $p(w_{1:D})$ of a document belonging ta a topic structure is intractable [blei2012]. Furthermore, LDA introduces priors, so that $\beta_{k}\thicksim Dirichlet(\eta)$ and $\theta_{d}\thicksim Dirichlet(\alpha)$.
@@
Though the topic structure cannot be efficiently computed, there are ways to approximate it [blei2003]. The two most prominent alternatives for this are Markov Chain Monte Carlo (MCMC) sampling methods and variational inference [jordan1999]. There are also online variants, such as online LDA [hoffman2010], where documents arrive in streams and $\lambda_{1:K}$ is updated according to each document batch.
@@
The work of [mamakos2018overlapping] was the first to employ PTM for multiagent learning, while existing literature on multiagent learning [fudenberg1998theory, tuyls2012multiagent], in both non-cooperative [hu2003nash, hu1998multiagent, littman1994markov] and cooperative [DBLP:conf/ijcai/BalcanPZ15, DBLP:conf/atal/ChalkiadakisB04, DBLP:conf/atal/KrausST03, DBLP:conf/atal/KrausST04] game settings, had been largely preoccupied with the study of RL, PAC learning, or simple belief updating algorithms. According to it, each agent $i$ (out of $n$) maintains her own model. For each coalition $C$ $\in C, ii$ observes the achieved utility $u_{C}$. The contribution $r_{i,C}$ of agent $i$ to coalition $C$ is known to each other agent $j\in C\setminus i$, once $C$ is successfully formed. To be able to feed this information to their models, agents maintain a vocabulary. The agent vocabulary includes $n$ words, one for each agent, indicating their contribution, plus two words for the utility, one representing the gains and the other representing the losses, summing up to $n+2$ words in total. Assuming a game round $t$, the agent $i$ interprets the coalition configurations of $C,i\in C$, as a document by appending in the document the word that indicates the contribution of agent $j\in C$$r_{j,C}$ times—where $r_{j,C}\in\mathbb{N}^{+}$ is the contribution of $j$ to $C$. Since words are discrete data, $u_{C}$ cannot be real-valued. To overcome this, we can apply a simple rounding operation to its value.
@@
Another, “non-conventional” use of PTM is given in [georgara2019learning]. In this setting, agents form coalitions that participate in an unknown hedonic game [aziz2016hedonic], and can observe only a few game instances. The document in this case corresponds to agent identifiers in a coalition and a binary label (‘gain’ or ‘loss’) indicating the sign of the revenue that accrues by this coalition. By applying online LDA it is possible to discover ordinal agent collaboration preferences and learn the latent game parameters.
@@
Finally, in another work, PTM has been combined with RL to learn user preferences and generate movie recommendations [tripolitakis2017probabilistic]. This approach models items and users as distributions of topics with Dirichlet priors, leveraging crowdsourced corpora e.g. Wikipedia movie synopses. Explicit attributes are substituted by latent features in both items and users distributions and consequently reinforcement learning is applied for recommendation policy adaptation. In RL, the learning rate is altered according to if the generated recommendations are close to, or far from the average recommendations that the agent has been giving, this way minimizing changes during strong performance and significantly increasing it to escape from low-rated recommendations.

### v3

@@ 3.1 Multiagent Reinforcement learning
Another notable paper is [30] that introduces a fully scalable and decentralized MADRL algorithm, the Multiagent Advantage Actor Critic (MA2C). MA2C is built as an extension to the well-known DRL algorithm Advantage Actor Critic (A2C) and is designed to tackle the adaptive traffic signal control problem. It does so, by enabling limited communication between the agents, and specifically sharing fingerprints of policies between neighbors. The connections between the agents help to alleviate the issues caused by partial observability and non-stationarity, as it helps each entity to have more informed decision-making based on the anticipated actions of its neighbors. Another key feature of this framework, is that it tailors the reward of each agent $i$, by computing a discounting, weighted, global reward with increased weights for the rewards of agents in agent’s $i$ neighborhood. Essentially, each individual’s reward is discounted according to the agent’s distance from agent $i$. MA2C is evaluated on both a large synthetic traffic grid and the real-world traffic network of Monaco city. In this context, the algorithm demonstrates superior performance of optimality, robustness and sample efficiency over traditional decentralized MADRL algorithms and other state-of-the-art methods.
@@
Despite the fact that they can be criticized as lacking in theoretical foundations in a multiagent setting, independent learners (IL) can attain high performance in hard multiagent problems. The underpinning of IL algorithms is that each agent learns on its own and perceives the rest of the agents as part of the environment with no communication in between, thus pertaining to the SIH. This simple logic can surprisingly attain performance comparable, and sometimes even better than more sophisticated techniques [34, 106, 156], and simultaneously be scalable due to its by-design decentralization. The first algorithm introduced in this context is a value-based algorithm based on $Q$-learning, the Independent $Q$-learners (IQL) [135]. According to IQL, each agent maximizes their own $Q$-function independently and optimize their own policy without sharing any information and without cooperating. A more modern approach adapts PPO in independent multiagent environments to present the actor-critic algorithm called Independent Proximal Policy Optimization (IPPO) [34]. In a similar fashion to IQL, IPPO does not require any information sharing or communication; instead, each agent works towards optimizing its own policy. Despite that, IPPO has achieved very good performance in hard problems such as the Starcraft Multiagent Challenge. Although IL can achieve competitive results and at the same time be straightforward and simplistic, it may struggle with the non-stationarity of the environment that is introduced by other agents’ behavior, since each decision-maker acknowledges its equivalents as dynamic parts of the environment.
@@
A relevant algorithm that uses a more sophisticated approach and is based on PPO, is that of Multiagent Proximal Policy Optimization (MAPPO) [156] that achieves good results in several cooperative benchmarks. MAPPO is an actor-critic algorithm, that follows a centralized training architecture which in turn leverages global information as the critic is trained on the joint state value function, making the agents more altruistic and disregarding the SIH. Additionally, there is no communication incorporated in this technique since the agents train in a centralized manner. The result is a sample-efficient algorithm that, unlike IL approaches, mitigates the non-stationarity problem making it less susceptible to changes in the environment made by other learners. As its single-agent counterpart, MAPPO serves as a very strong baseline due to its high performance in most cooperative environments and its robustness to hyperparameter tuning [106].
@@
Another interesting approach is the Value Decomposition Networks (VDN) [132]. VDN is a groundbreaking algorithm that is designed to train multiple agents to work together towards achieving a common goal. Its key point, as the name suggests, is the principle of decomposing the joint value function into a sum of individual agents’ value functions. Ultimately, each agent builds a policy that maximizes the sum of policies which serves as a proxy for the joint value function. This offers a subtle way of coordination without the use of explicit communication mechanisms, since each entity makes decisions based on its perspective while indirectly contributing to collective intelligence. By utilizing the sum of these individual contributions and matching local rewards to the corresponding agents, VDN tackles two inherent MAS pathogenies such as the credit assignment problem and scalability. This method has opened up new avenues in research for designing sophisticated MAS capable of complex interactions and deeply strategic behavior.
@@
Finally, the Cooperative Multi-Agent Exploration (CMAE) [79] is an exploration strategy that works on top of any MARL algorithm and promotes coordination through shared goals, particularly in sparse-reward environments. CMAE innovates by projecting the complex, high-dimensional state space into simpler, lower-dimensional “restricted spaces” for more efficient exploration. Using a hierarchical approach, exploration starts from these simpler spaces and expands as needed. This mechanism reshapes the rewards within the replay buffer to incentivize agents towards common objectives. It relies on entropy-based techniques for goal selection and demonstrates significant improvements in learning efficiency and effectiveness in tasks within environments like the Multiple Particle Environment and the StarCraft Multi-Agent Challenge, showcasing its superiority over traditional exploration methods.
@@
MA2C [30] ATOC [63] VDN [132] MAPPO [156] IQL [135] IPPO [34]
@@ 4 Rationality Solution Concepts
In complex multi-agent domains, understanding Nash equilibrium (NE) and other rationality solution concepts is vital for designing agents that interact effectively, as well as analyzing and providing insights on their behavior and outcomes. Usually, the desired properties of the outcomes in this scope are fairness and stability. Fairness corresponds to how well the agents’ payoffs comply with their actual contributions, while stability reflects the non-willingness of agents to select a deviating strategy [24]. In this section, we will present and discuss such game theoretic notions within the field of multi-agent DRL, with an emphasis on correlated equilibrium (CE), and on fairness-related solution concepts.
To begin with, the well-known Nash theorem guarantees that every finite two-player game has at least one NE. In games where multiple equilibria exist, one has to search for the most desirable one among them. One measure of desirability is the degree of social welfare maximization, i.e. the sum of all players’ payoffs. The search for an NE is PPAD-complete, however finding an NE with maximal social welfare is NP-hard even in symmetric 2-player games, making it computationally challenging. There have been attempts to reach approximate Nash equilibria that achieve near optimal social welfare. In [31] it is shown that this can be discovered in polynomial time.
@@
Moreover, there exist certain successful integrations of CE in multi-agent DRL. As an example, in [157] the authors apply CE on smart generation control in power systems. They present a novel algorithm that learns to achieve optimal joint equilibrium strategies. This approach employs the correlated equilibria in order to facilitate optimal coordination and consequently improve performance. Similarly, in [42] they employ a dynamic correlation matrix to model agent interactions and complex relationships. With this technique, the authors enable the discovery of optimal CE and enhance collective outcomes. That work highlights how valuable are CE in hostile environments in which agent interconnections are complex and change dynamically. Another successful application is presented in [55], where multiple agents employ the DQN algorithm to achieve CE optimal solutions in order to optimize a textile manufacturing optimization process.
In [47], authors employ DRL for training agents that participate into uniform price auctions. Interestingly, this Deep Deterministic Policy Gradient (DDPG) algorithm leads to learned policies which are shown to converge to NE. Also, this algorithm can tackle continuous state and action spaces, a feature that Q-learning does not possess, though a drawback of this approach is that it requires fine tuning of the related parameters.
@@
A work that addresses fair solutions is [63]. In that work, authors present a novel Fair-Efficient Network (FEN) model that manages to learn both efficiency and fairness in multiagent domains. This is achieved via fairness decomposition into objectives and utilizes a hierarchical structure, which contains a controller and multiple sub-entities, each one maintaining a distinct policy. This method supports fairness and addresses the challenges of stability in learning environments, as it can scale to larger problems due to the model’s ability to facilitate decentralized learning. Another notable contribution, is that of [165], where a metric is put forward that evaluates fairness in the form of optimizable welfare functions. A novel NN architecture is used that consists of sub-networks and is capable of optimizing the different aspects of fairness such as efficiency and equity in the form of welfare functions. Ultimately, this technique leads to fair outcomes. Additionally, in [144], a novel framework for controlling Unmanned Aerial Vehicles (UAVs) and optimizing their trajectories in terms of both performance and fairness is presented. In this context, authors propose a method where each UAV maximizes its reward independently and, at the same time, a collective approach for maximizing fairness is maintained.
@@
Finally, we list the most common solution concepts that can be found in the MARL literature. These are the Shapley value [120], the Banzhaf index [11], the core [44], the nucleolus [116], the kernel [33], the bargaining set [6], and the stable set [142].
The Shapley value guarantees that an agent will receive a payoff that is proportional to its contribution. Considering a characteristic function game $G=(N,v)$ with $N$ agents and $v$ the characteristic function, then the Shapley value $\phi_{i}(G)$ is calculated based on each agent’s $i$ marginal contribution:
@@
While the Shapley value and the Banzhaf index focus on the fairness of the payoffs, there are also solution concepts that deal with the stability of the formed coalitions in the game. The core solution concept is the set of all outcomes $(CS,\mathbf{x})$, where $CS$ is a coalition structure and $\mathbf{x}$ the payoff vector, for which $x(C)\geq v(C)$ for every $C\subseteq N$, where $x(C)$ denotes the total payoff for $C$ under $\mathbf{x}$. Note that not every game has a non-empty core.
@@
The works of [26, 25] combined the concept of Bayesian core, a cooperative solution concept, with non-cooperative solution concepts. This is inline with the our objectives, since we envision to revolutionize the way we think about equilibria, and pave the road towards the discovery of novel, meaningful, and practical concepts of rationality, so as to enable computationally tractable strategic deliberations in uncertain multiagent environments. Recall that a grand challenge put forward early in this paper, is that of overriding restricting GT assumptions, such as the SIH and CPA, [^1] Note that the Bayesian core solution concept does not necessitate the use of the CPA assumption. and to identify tractable ways for computing initial heterogeneous agent beliefs into the focus. This could potentially give rise to novel, practical equilibrium concepts, which explicitly take the initial agent beliefs-forming stage into account - and which come complete with bounded rationality capabilities and anytime performance guarantees. Such equilibrium concepts would have the potential to be employed in cooperative and non-cooperative domains alike, rendering as such the distinction between cooperative and non-cooperative GT solution concepts artificial and obsolete.
@@ 5 Probabilistic Topic Modeling
Existing Probabilistic Topic Modeling (PTM) algorithms have been applied successfully in various domains, such as document and text classification, image analysis and organization, movie recommendation systems, and population genetics [15]. One could potentially adapt existing PTM algorithms for opponent modeling or strategy modeling purposes in multiagent environments—i.e., use them to derive the prior beliefs of the agents. Note that the ability to assign meaningful probability to previously unseen “documents” is inherent in many existing PTM algorithms (such as the LDA). Note also that PTMs constitute Bayesian exploration techniques, assigning as they do “topics” to modeled items according to probabilistic beliefs embodied in the mixture distributions they maintain. As such, they do not require the use of explicit heuristics or stoppage conditions to tackle the exploration-exploitation trade-off that is ever-present in decision-making problems.
Probabilistic topic models (PTMs) utilize statistical methods to analyze words in documents, so as to discover the underlying topics and their interconnection. The most popular implementation is based on the Latent Dirichlet Allocation (LDA) [14].
@@
However, computing the probability $p(w_{1:D})$ of a document belonging ta a topic structure is intractable [15]. Furthermore, LDA introduces priors, so that $\beta_{k}\thicksim Dirichlet(\eta)$ and $\theta_{d}\thicksim Dirichlet(\alpha)$.
@@
Though the topic structure cannot be efficiently computed, there are ways to approximate it [14]. The two most prominent alternatives for this are Markov Chain Monte Carlo (MCMC) sampling methods and variational inference [64]. There are also online variants, such as online LDA [56], where documents arrive in streams and $\lambda_{1:K}$ is updated according to each document batch.
@@
The work of [87] was the first to employ PTM for multiagent learning, while existing literature on multiagent learning [41, 138], in both non-cooperative [59, 58, 77] and cooperative [10, 23, 72, 73] game settings, had been largely preoccupied with the study of RL, PAC learning, or simple belief updating algorithms. According to it, each agent $i$ (out of $n$) maintains her own model. For each coalition $C$ $\in C, ii$ observes the achieved utility $u_{C}$. The contribution $r_{i,C}$ of agent $i$ to coalition $C$ is known to each other agent $j\in C\setminus i$, once $C$ is successfully formed. To be able to feed this information to their models, agents maintain a vocabulary. The agent vocabulary includes $n$ words, one for each agent, indicating their contribution, plus two words for the utility, one representing the gains and the other representing the losses, summing up to $n+2$ words in total. Assuming a game round $t$, the agent $i$ interprets the coalition configurations of $C,i\in C$, as a document by appending in the document the word that indicates the contribution of agent $j\in C$$r_{j,C}$ times—where $r_{j,C}\in\mathbb{N}^{+}$ is the contribution of $j$ to $C$. Since words are discrete data, $u_{C}$ cannot be real-valued. To overcome this, we can apply a simple rounding operation to its value.
@@
Another, “non-conventional” use of PTM is given in [43]. In this setting, agents form coalitions that participate in an unknown hedonic game [8], and can observe only a few game instances. The document in this case corresponds to agent identifiers in a coalition and a binary label (‘gain’ or ‘loss’) indicating the sign of the revenue that accrues by this coalition. By applying online LDA it is possible to discover ordinal agent collaboration preferences and learn the latent game parameters.
@@
Finally, in [136], PTM has been combined with RL to learn user preferences and generate movie recommendations. This approach models items and users as distributions of topics with Dirichlet priors, leveraging crowdsourced corpora e.g. Wikipedia movie synopses. Explicit attributes are substituted by latent features in both items and users distributions and consequently reinforcement learning is applied for recommendation policy adaptation. In RL, the learning rate is altered according to if the generated recommendations are close to, or far from the average recommendations that the agent has been giving, this way minimizing changes during strong performance and significantly increasing it to escape from low-rated recommendations.